OpenStudy (anonymous):

An amateur drama company were selling tickets for their latest production. the opening night was a sell out with all 200 tickets being sold. Tickets were priced as follows: Adults $9.50 each Children $4.50 each Senior citizens $3.50 each The takings for the opening night amounted to $1375. Also, there were three times as many children's tickets sold as there were senior citizens' tickets. a) construct a system of linear simultaneous equations to represent this purchase of tickets for the opening night. Determine algebraically the number of each type of ticket sold for the opening night.

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

I'm having trouble with constructing the linear simultaneous equations. one of them will be this I believe but I'm unsure about the rest. a=adults, c=children, s=senior 9.50a+4.50c+3.50s =1375

OpenStudy (mertsj):

And c = 3s

OpenStudy (mertsj):

And a + c + s = 200

OpenStudy (mertsj):

So in yours two equations, replace c with 3s and solve the system.

OpenStudy (anonymous):

kk. let me try it. I also got a + c + s = 200 but then I was unsure about the c=3s because I thought it was a + 3s + s = 200, it's really confusing for me.

OpenStudy (mertsj):

Right. It is. So a + 4s = 200

OpenStudy (anonymous):

yes, I solved it correctly. I just had trouble with figuring out the equations, thanks so much!

OpenStudy (mertsj):

yw